OpenStudy (anonymous):

Given that events A and B are independent, what is P (A|B)? P (A) P (B) P (A and B) / P (B) P (A or B) P (A and B)

myininaya (myininaya):

\[P(A|B)=\frac{P(A and B)}{P(B)}\] for any events A and B BUT!!!! this can be simplified further since A and B are independent

myininaya (myininaya):

P(A and B)= to what if A and B are independent

OpenStudy (anonymous):

so its the last answer? P(A and B)

myininaya (myininaya):

I didn't say that

myininaya (myininaya):

I'm asking you to tell me what P(A and B) equals given that A and B are independent

OpenStudy (anonymous):

im confused

myininaya (myininaya):

here are your hints: \[\text{ for any events } A \text{ and } B \text{ we have } P(A|B)=\frac{P(A \text{ and } B)}{P(B)} \\ \text{ if } A \text{ and } B \text{ are independent then } P(A \text{ and } B)=P(A) \cdot P(B)\]

myininaya (myininaya):

use these two things to get your simplified answer for what is P(A|B) if A and B are independent

OpenStudy (anonymous):

oh so it would be just P(A)

myininaya (myininaya):

yes but technically P(A and B)/P(B) is also right it is just not completely simplified but if I only got once choice I would go with P(A)

OpenStudy (anonymous):

are you sure :/

OpenStudy (anonymous):

yes sure

OpenStudy (anonymous):

satellite, do you agree with P(A)

OpenStudy (misty1212):

it is the definition of independent that A and B are INDEPENDENT if \[P(A|B)=P(A)\] that is exactly what the word means

OpenStudy (anonymous):

i agree with misty